# The Panaceas for Improving Low-Rank Decomposition in Communication-Efficient Federated Learning

### v1

@@ Appendix B Discussion about the Upper Rank Bound of BKD
Appendix C Proof of Theorem 1
In this section, we first present the problem formulation, then outline the assumptions and lemmas necessary for the proof, and conclude with a detailed proof process. Our theoretical analysis relies mainly on FedLMT (Liu et al., 2024).
C.1 Problem Formulation
@@
$\mathbf{x}_{i}^{t}=\left\{\begin{array}[]{lr}\mathbf{x}_{i}^{t-1}-\eta\frac{1}{N}\sum_{j=1}^{N}\nabla H_{j}(\mathbf{x}_{j}^{t-1},\xi_{j}^{t})&\text{if}\ t\mod E=0,\\ \mathbf{x}_{i}^{t-1}-\eta\nabla H_{i}(\mathbf{x}_{i}^{t-1},\xi_{i}^{t})&\text{otherwise},\\ \end{array}\right.$ (15)
@@
$\mathbf{x}_{i}^{t}=\left\{\begin{array}[]{lr}\text{reset\_updates}(\mathbf{x}_{i}^{t})&\text{if}\ t\mod S=0,\\ \mathbf{x}_{i}^{t}&\text{otherwise}.\\ \end{array}\right.$ (16)
@@
$\nabla h_{i}(\mathbf{x}_{i}^{t-1})$ =$=\{\nabla h_{i}^{1}(\mathbf{x}_{i}^{t-1}),\dots,\nabla h_{i}^{L}(\mathbf{x}_{i}^{t-1})\},\quad\nabla H_{i}(\mathbf{x}_{i}^{t-1})=\{\nabla H_{i}^{1}(\mathbf{x}_{i}^{t-1}),\dots,\nabla H_{i}^{L}(\mathbf{x}_{i}^{t-1})\},$ (23)
$\nabla f_{i}(\mathbf{w}_{i}^{t-1})$ =$=\{\nabla f_{i}^{1}(\mathbf{w}_{i}^{t-1}),\dots,\nabla f_{i}^{L}(\mathbf{w}_{i}^{t-1})\},\quad\nabla F_{i}(\mathbf{w}_{i}^{t-1})=\{\nabla F_{i}^{1}(\mathbf{w}_{i}^{t-1}),\dots,\nabla F_{i}^{L}(\mathbf{w}_{i}^{t-1})\},$
@@ C.3 Lemmas
$\mathbb{E}\left[\left\|\bar{\mathbf{w}}^{t}-\mathbf{w}_{i}^{t}\right\|_{2}^{2}\right]\leq\Lambda\triangleq 4\rho\eta^{2}E^{2}G^{2}+12(L-\rho)\eta^{2}E^{2}G^{2}(\Gamma_{u}^{4}+\Gamma_{v}^{4}+\eta^{2}E^{2}G^{2}\Gamma_{u}^{2}\Gamma_{v}^{2}).$ (38)
@@
$=$ $3\sum_{l=\rho+1}^{L}(4\eta^{2}E^{2}G^{2}\Gamma_{u}^{4}+4\eta^{2}E^{2}G^{2}\Gamma_{v}^{4}+4\eta^{4}E^{4}G^{4}\Gamma_{u}^{2}\Gamma_{v}^{2})=12(L-\rho)\eta^{2}E^{2}G^{2}(\Gamma_{u}^{4}+\Gamma_{v}^{4}+\eta^{2}E^{2}G^{2}\Gamma_{u}^{2}\Gamma_{v}^{2})$
@@
=$\overset{(a)}{=}$ $\sum_{l=1}^{\rho}\eta^{2}\mathbb{E}\left[\left\|\frac{1}{N}\sum_{i=1}^{N}\left[\nabla F_{i,l}(\mathbf{w}_{i}^{t-1})-\nabla f_{i,l}(\mathbf{w}_{i}^{t-1})\right]\right\|_{F}^{2}\right]+\sum_{l=1}^{\rho}\eta^{2}\mathbb{E}\left[\left\|\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i,l}(\mathbf{w}_{i}^{t-1})\right\|_{F}^{2}\right]$
@@
=$\overset{(a)}{=}$ $\sum_{l=1}^{\rho}(-\eta)\mathbb{E}\left[\mathbb{V}(\nabla f_{l}(\bar{\mathbf{w}}^{t-1}))^{\top}\cdot\mathbb{V}(\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i,l}(\mathbf{w}_{i}^{t-1}))\right]$
=$\overset{(b)}{=}$ $\sum_{l=1}^{\rho}\frac{-\eta}{2}\mathbb{E}\left[\left\|\mathbb{V}(\nabla f_{l}(\bar{\mathbf{w}}^{t-1}))\right\|_{2}^{2}+\left\|\mathbb{V}(\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i,l}(\mathbf{w}_{i}^{t-1}))\right\|_{2}^{2}\right]$
@@
=$\overset{(c)}{=}$ $\sum_{l=1}^{\rho}\frac{-\eta}{2}\mathbb{E}\left[\left\|\nabla f_{l}(\bar{\mathbf{w}}^{t-1})\right\|_{F}^{2}+\left\|\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i,l}(\mathbf{w}_{i}^{t-1})\right\|_{F}^{2}\right]$
@@
where (a) follows from $f_{i}(\mathbf{w})\triangleq\mathbb{E}_{\xi_{i}\in\mathbb{D}_{i}}[F_{i}(\mathbf{w},\xi_{i})]$. (b) follows from $-\mathbf{a}^{\top}\mathbf{b}=\frac{1}{2}\left[\left\|\mathbf{a}-\mathbf{b}\right\|_{2}^{2}-\left\|\mathbf{a}\right\|_{2}^{2}-\left\|\mathbf{b}\right\|_{2}^{2}\right]$ for two column vectors $\mathbf{a}$ and $\mathbf{b}$. (c) follows from $\left\|\mathbb{V}(X)\right\|_{2}^{2}=\left\|X\right\|_{F}^{2}$ for any matrix $X$. (d) follows from Eq.(LABEL:eq:use_Lambda).
@@
=$\overset{(a)}{=}$ $\sum_{l=\rho+1}^{L}(-\eta)\underbrace{\mathbb{E}\left[\mathbb{V}(\nabla f_{l}(\bar{\mathbf{w}}^{t-1})\bar{V}_{l}^{t-1})^{\top}\cdot\mathbb{V}(\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i,l}(\mathbf{w}_{i}^{t-1})V_{i,l}^{t-1})\right]}_{A_{l}}$
@@ C.4 Proof of Theorem 1
Under Assumptions 1, 2, 3 and 4, let $21<c<$ 2 be a constant and the learning rate satisfy $0<\eta\leq\min\{(\frac{\psi_{uv}^{2}}{2})^{\frac{1}{c-1}},\frac{1}{L_{s}},1\}$, we have
@@
where (a) follows from $0<\eta\leq 1$ and $21<c<$ 2, thus $-\eta\leq-\eta^{c}$ and $\eta^{2}\leq\eta^{c}$. (b) follows from $\eta\leq(\frac{\psi_{uv}^{2}}{2})^{\frac{1}{c-1}}$, and thus $2\eta^{c}-\eta\psi_{uv}^{2}\leq 0$. (c) follows from $\eta\leq\frac{1}{L_{s}}$, and thus $L_{s}\eta^{2}-\eta\leq 0$. Therefore, we have

### v2

@@ Appendix B Discussion about the Upper Rank Bound of BKD
Appendix C Connection to More Fields
Our work focuses on communication compression in federated learning, which can be regarded as a specific form of model compression. In deep learning, a variety of model compression techniques have been developed to reduce computational and storage costs while preserving model performance (Li et al., 2023, 2024c, 2024b). Beyond efficiency, recent research underscores the growing importance of neural network interpretability (Liu et al., 2025a, b, 2024c) and causal reasoning (Liu et al., 2023b, 2024b, a). Building on these trends, we view interpretable communication compression as a promising avenue for future investigation.
Appendix D Proof of Theorem 1
In this section, we first present the problem formulation, then outline the assumptions and lemmas necessary for the proof, and conclude with a detailed proof process. Our theoretical analysis relies mainly on FedLMT (Liu et al., 2024a).
D.1 Problem Formulation
@@
[…]
@@
[…]
@@
$\nabla h_{i}(\mathbf{x}_{i}^{t-1})$ $=\{\nabla h_{i}^{1}(\mathbf{x}_{i}^{t-1}),\dots,\nabla h_{i}^{L}(\mathbf{x}_{i}^{t-1})\},\quad\nabla H_{i}(\mathbf{x}_{i}^{t-1})=\{\nabla H_{i}^{1}(\mathbf{x}_{i}^{t-1}),\dots,\nabla H_{i}^{L}(\mathbf{x}_{i}^{t-1})\},$ (23)
$\nabla f_{i}(\mathbf{w}_{i}^{t-1})$ $=\{\nabla f_{i}^{1}(\mathbf{w}_{i}^{t-1}),\dots,\nabla f_{i}^{L}(\mathbf{w}_{i}^{t-1})\},\quad\nabla F_{i}(\mathbf{w}_{i}^{t-1})=\{\nabla F_{i}^{1}(\mathbf{w}_{i}^{t-1}),\dots,\nabla F_{i}^{L}(\mathbf{w}_{i}^{t-1})\},$
@@ D.3 Lemmas
[…]
@@
[…]
@@
$\overset{(a)}{=}$ $\sum_{l=1}^{\rho}\eta^{2}\mathbb{E}\left[\left\|\frac{1}{N}\sum_{i=1}^{N}\left[\nabla F_{i,l}(\mathbf{w}_{i}^{t-1})-\nabla f_{i,l}(\mathbf{w}_{i}^{t-1})\right]\right\|_{F}^{2}\right]+\sum_{l=1}^{\rho}\eta^{2}\mathbb{E}\left[\left\|\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i,l}(\mathbf{w}_{i}^{t-1})\right\|_{F}^{2}\right]$
@@
$\overset{(a)}{=}$ $\sum_{l=1}^{\rho}(-\eta)\mathbb{E}\left[\mathbb{V}(\nabla f_{l}(\bar{\mathbf{w}}^{t-1}))^{\top}\cdot\mathbb{V}(\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i,l}(\mathbf{w}_{i}^{t-1}))\right]$
$\overset{(b)}{=}$ $\sum_{l=1}^{\rho}\frac{-\eta}{2}\mathbb{E}\left[\left\|\mathbb{V}(\nabla f_{l}(\bar{\mathbf{w}}^{t-1}))\right\|_{2}^{2}+\left\|\mathbb{V}(\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i,l}(\mathbf{w}_{i}^{t-1}))\right\|_{2}^{2}\right]$
@@
$\overset{(c)}{=}$ $\sum_{l=1}^{\rho}\frac{-\eta}{2}\mathbb{E}\left[\left\|\nabla f_{l}(\bar{\mathbf{w}}^{t-1})\right\|_{F}^{2}+\left\|\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i,l}(\mathbf{w}_{i}^{t-1})\right\|_{F}^{2}\right]$
@@
where (a) follows from $f_{i}(\mathbf{w})\triangleq\mathbb{E}_{\xi_{i}\in\mathbb{D}_{i}}[F_{i}(\mathbf{w},\xi_{i})]$. (b) follows from $\mathbf{a}^{\top}\mathbf{b}=\frac{1}{2}\left[\left\|\mathbf{a}-\mathbf{b}\right\|_{2}^{2}-\left\|\mathbf{a}\right\|_{2}^{2}-\left\|\mathbf{b}\right\|_{2}^{2}\right]$ for two column vectors $\mathbf{a}$ and $\mathbf{b}$. (c) follows from $\left\|\mathbb{V}(X)\right\|_{2}^{2}=\left\|X\right\|_{F}^{2}$ for any matrix $X$. (d) follows from Eq.(LABEL:eq:use_Lambda).
@@
$\overset{(a)}{=}$ $\sum_{l=\rho+1}^{L}(-\eta)\underbrace{\mathbb{E}\left[\mathbb{V}(\nabla f_{l}(\bar{\mathbf{w}}^{t-1})\bar{V}_{l}^{t-1})^{\top}\cdot\mathbb{V}(\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i,l}(\mathbf{w}_{i}^{t-1})V_{i,l}^{t-1})\right]}_{A_{l}}$
@@ D.4 Proof of Theorem 1
Under Assumptions 1, 2, 3 and 4, let $1<c<2$ be a constant and the learning rate satisfy $0<\eta\leq\min\{(\frac{\psi_{uv}^{2}}{2})^{\frac{1}{c-1}},\frac{1}{L_{s}},1\}$, we have
@@
where (a) follows from $0<\eta\leq 1$ and $1<c<2$, thus $\eta\leq-\eta^{c}$ and $\eta^{2}\leq\eta^{c}$. (b) follows from $\eta\leq(\frac{\psi_{uv}^{2}}{2})^{\frac{1}{c-1}}$, and thus $2\eta^{c}-\eta\psi_{uv}^{2}\leq 0$. (c) follows from $\eta\leq\frac{1}{L_{s}}$, and thus $L_{s}\eta^{2}-\eta\leq 0$. Therefore, we have